HUBBLE SPACE TELESCOPE DAILY REPORT #824 PERIOD COVERED: 0800 EST Daily Status Report as of 047/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 4766 (Boron in Population Two Dwarf Stars, Part 2) The High Resolution Spectrograph was used to make observations of the Boron lines at 2496 and 2497 A0 in the metal poor unevolved star HD140283. The OSS monitored the spiral search portion of the onboard target acquisition. The target was found in the first map. No other problems were reported. 1.2 Completed FOC/96 3325 (Structure of Globular Clusters) The Faint Object Camera was used to make observations of the inner core regions of the globular cluster Omega Centauri (aka,NGC-5139). Wide Field Camera images were also taken in parallel with the FOC f/96 camera. The observations were made during LOS, and still remain of the science tape recorder. They will be analyzed later. 1.3 Completed FOC/96 4353 (Far UV Observations of the Giant Planets) The Faint Object Camera (f/96)was used to make ultraviolet observations of the northern polar regions of the planet Jupiter. The first observations was received in real time and displayed for the principal investigator. A corrective SIP maneuver was successfully executed, and the following observation occurred during LOS. The Play back showed significant jitter, although FGS #3 remained in course track. During the next observation an "interactive acquisition" was scheduled, but the scheduled corrective SIP maneuvers were canceled because the pointing correction was successfully uplinked at the first opportunity. The science observations still remain on the Science Tape Recorder, and will be analyzed later. No problems were reported. 1.4 Completed FOS/BL 4698 (FOS Dark Monitoring) The Faint Object Spectrograph (Blue Side) was used to make dark calibration exposures to determine the instruments background. The observations were executed as scheduled, and monitored in the OSS. No problems were reported. 1.5 Initiated FOC/96 4071 (UV Studies of Galaxies) The Faint Object Camera (f/96) was used to make observations of the structure of the nuclei of galaxy NGC-4278. This is to study the Oxygen II emission regions. The observations were monitored in the OSS, and no problems were reported. 1.6 Initiated WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to take high latitude observations in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 11 Successful : 11 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Significant HSTARs: None 2.4 Operations Note: The FGS #2 anomaly is still under investigation and an analysis review meeting will be held this afternoon (Tuesday, Feb. 16). 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 4071 (UV Studies of Galaxies) 3.2 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.3 Complete WFC/PC 4109 (Rapid Internal Monitor) 3.4 Complete Two Sets of FOC/96 3591 (Mass Stars in Star Burster Galaxies, FOC Images, Part 1) 3.5 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.6 Complete GHRS/2 4792 (Hydrogen Lyman Alpha Day Glow Emission Line Profiles) 3.7 Complete FOS/RD 3791 (Quasar Absorption Line Survey) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END- ----- End Included Message -----